alice and dan's ages add up to 110. dan is two years older than alice. how old are alice and dan
Answer by only4christ(12)

Let A = Alice's age and D = Dan's age
Then, A + D = 110 and D = A + 2
Using the substitution method, we will substitute the equation D=A+2 into the first equation to get:
A + A + 2 = 110
2A + 2 = 110 Collect like terms
2A = 108 move the 2 to the other side of the equation
A = 54, which is Alice's age
D = A + 2 so D = 54 + 2 = 56, which is Dan's age